OpenStudy (anonymous):

6x4 + 5x3 + 3x - 5 is divided by 3x2 - 2x in synthetic form

OpenStudy (anonymous):

24 +15+3x-5 => 3x+34 divided by 6-2x. as we put 6-2x=0 => 2(3-x)=0 3-x=0=>x=3 now by the synthetic division method.... as the equation has coefficients.... 3 34 so 3)3 34(1+11 3/ 0 34 33/1 (here 1 is remainder ) so 1x+11x (by starting power from right side) then answer wll be x+11.

OpenStudy (anonymous):

thanks .. i need help with 5x3 + 6x + 8 by x + 2

OpenStudy (anonymous):

here 15+6x+8 6x+23 by x+2

OpenStudy (anonymous):

ok